OpenStudy (goformit100):

Which of the following is not a way to represent the solution of the inequality 3(2x - 1) greater than or equal to 4(2x - 3) - 3? x less than or greater to 6 6 greater than or equal to x A number line with a closed circle on 6 and shading to the left. A number line with a closed circle on 6 and shading to the right.

OpenStudy (goformit100):

@texaschic101

OpenStudy (texaschic101):

3(2x - 1) >= 4(2x - 3) - 3 6x - 3 >= 8x - 12 - 3 6x - 8x >= -15 + 3 -2x >= - 12 (change sign when dividing by a negative) x <= 6 Incorrect one is the last answer choice